# Numerical Hopf–Lax formulae for Hamilton–Jacobi equations on unstructured geometries

### v1

@@ 3 Construction of the scheme, stationary case
[…]
where $\mathcal{A}$ stands for the set of measurable functions mapping $[0,+\infty]$ into $\mathbb{R}^{d}$, and $y(s;\alpha)$ solves the ODE system
$\begin{cases}\dot{y}(s;\alpha)=\alpha(s),\quad s>0\\ y(0,\alpha)=x.\end{cases}$
In view of the $x$-dependence of $f$, (16) cannot be given by an exact Hopf–Lax type form; therefore, setting $y=y(\Delta t;\alpha)$, and $\alpha(s)\equiv(x-y)/\Delta t$, and assuming smoothness as usual, we approximate the integral in (16) as
[…]
@@
$v_{j}=\min_{k\in\mathbb{N}}\left\{e^{-\lambda\Delta t}v_{k}+\Delta tH^{*}\left(\frac{x_{j}-x_{k}}{\Delta t}\right)\right\}+\Delta tf(x_{j}),$ (19)
with obvious modifications in the case of Dirichlet boundary conditions.
Equation (19) is a fixed-point equation, and it is easy to show that it has a contractive right-hand side (see [7]); this ensures that it has a unique solution $V$. As for the convergence of $V$ to the solution of (7), it is possible to retrace the arguments in [7], along with those of the time-dependent case above, to obtain a consistency estimate in the form
@@
While approximating the integral as in (17) suffices to obtain a consistent scheme for $\Delta x=o(\Delta t)$, the term $O(\Delta t)$ becomes a bottleneck for the consistency rate. An easy computation shows that the truncation error (20) is maximized taking $\Delta t\sim\Delta x^{2/3}$, and accordingly the consistency rate w.r.t. $\Delta x$ drops to 2/3. A way to increase the rate is to approximate the integral of $f$ by a trapezoidal quadrature rule,
[…]
which results in the scheme
$v_{j}=\min_{k\in\mathbb{N}}\left\{e^{-\lambda\Delta t}v_{k}+\Delta tH^{*}\left(\frac{x_{j}-x_{k}}{\Delta t}\right)+\frac{\Delta t}{2}e^{-\lambda\Delta t}f(x_{k})\right\}+\frac{\Delta t}{2}f(x_{j}).$ (22)
@@ 3.1 Value iteration
In the easiest approach, (19) may be solved by the so-called value iteration, that is, via the iteration
$v_{j}^{(n+1)}=\min_{k\in\mathbb{N}}\left\{e^{-\lambda\Delta t}v_{k}^{(n)}+\Delta tH^{*}\left(\frac{x_{j}-x_{k}}{\Delta t}\right)\right\}+\Delta tf(x_{j}),$ (24)
where the index $n$ should now be understood as an iteration number (an obvious modification applies for treating (22)). Since the right-hand side of (19) is a contraction, the iteration (24) converges to its unique fixed point. The contraction coefficient, however, is given by $e^{-\lambda\Delta t}$, and the convergence becomes very slow as the approximation is refined. This backs the use of faster solvers for the fixed-point equation, in particular the so-called Policy Iteration (PI), which will be briefly reviewed here.
@@ 3.2 Exact policy iteration
    $\alpha^{(n+1)}=\arg\min_{\alpha\in U^{N}}\left(B(\alpha)V^{(n)}-c(\alpha)\right),$
@@
$\min_{k\in\mathbb{N}}\left\{-v_{j}+e^{-\lambda\Delta t}v_{k}+\Delta t\left(H^{*}\left(\frac{x_{j}-x_{k}}{\Delta t}\right)+f(x_{j})\right)\right\}=0.$ (26)
To interpret (26) in the form (25), it suffices to define the elements of the matrix $B(k)$ and the vector $c(k)$ as
$b_{jm}(k)=\begin{cases}1&\text{ if }m=j\neq k\\ -e^{-\lambda\Delta t}&\text{ if }m=k\neq j\\ 1-e^{-\lambda\Delta t}&\text{ if }m=j=k\\ 0&\text{ otherwise};\end{cases}\hskip 42.67912ptc_{j}(k)=\Delta t\left(H^{*}\left(\frac{x_{j}-x_{k}}{\Delta t}\right)+f(x_{j})\right).$ (27)
In this case, denoting by $k_{j}$ the node index achieving the minimum in (26), and changing all the signs, the policy evaluation phase is rewritten as the system
$v_{j}-e^{-\lambda\Delta t}v_{k_{j}}=\Delta t\left(H^{*}\left(\frac{x_{j}-x_{k_{j}}}{\Delta t}\right)+f(x_{j})\right)\quad(j=1,\ldots,N).$ (28)
With these definitions, assumptions (P1)(P2) of Theorem 3 are satisfied. In fact, the continuity of $B$ and $c$ is obvious once the set of node indices is naturally endowed with the discrete topology. As for (P2), we note that an equivalent definition of monotone matrix requires the off-diagonal elements to be nonpositive (which is clearly the case), and the eigenvalues to have nonnegative real part; this latter property is true, as a consequence of Gershgorin’s theorem. Note that monotonicity of the matrices $B$ implies their nonsingularity, and thus the unique solvability of the policy evaluation phase.
Clearly, the form (22) of the scheme can be treated by changing the definition of the vector $c(k)$ into
$c_{j}(k)=\Delta t\left(H^{*}\left(\frac{x_{j}-x_{k}}{\Delta t}\right)+\frac{f(x_{j})+e^{-\lambda\Delta t}f(x_{k})}{2}\right),$
@@ 3.3 Modified policy iteration
An adaptation of policy iteration to large sparse problems has been proposed as “modified policy iteration” (MPI) in [9], and has also become a classical tool. In this case, the policy evaluation phase is carried out in iterative form, which may be convenient in very large problems, whenever a direct solver would be unfeasible. In the specific case of (26), once given an initial guess for the policy, the two phases of MPI can be written as
@@
    $v_{j,m+1}^{(n)}=e^{-\lambda\Delta t}v_{k_{j}^{(n)},m}+\Delta t\left(H^{*}\left(\frac{x_{j}-x_{k_{j}^{(n)}}}{\Delta t}\right)+f(x_{j})\right)$
@@
    $k_{j}^{(n+1)}=\arg\min_{k\in\mathbb{N}}\left\{-v_{j}^{(n)}+e^{-\lambda\Delta t}v_{k}^{(n)}+\Delta t\left(H^{*}\left(\frac{x_{j}-x_{k}}{\Delta t}\right)+f(x_{j})\right)\right\}.$
Here, we have denoted by $v_{j,m}^{(n)}$ the $j$-th element of the $m$-th iterate in the policy evaluation, by $v_{j}^{(n)}$ the corresponding final value of the iteration (i.e., the $j$-th element of $V^{(n)}$), and by $k_{j}^{(n)}$ the $j$-th element of the optimal policy at the $n$-th iteration.
Remark 2
Since the policy is defined, in this scheme, by the set of indices $k_{j}$ of the nodes achieving the minima for $\mathcal{F}_{j}$, we can consider the policy to be convergent as soon as none of these indices is any longer updated.
@@ 4 Unstructured implementation
We provide some details for the actual implementation of the fully discrete scheme under consideration, on two-dimensional triangular grids. The aim is to compute the global minimum in (9) as efficiently as possible, in terms of both space and time complexity, by means of walks along the grid nodes that locally decrease the right hand side of our scheme. To this end, once a triangulation (typically, Delaunay) is given, we only assume to work with the minimal data structures describing the domain geometry, namely vertex and triangle lists, plus a connectivity list which contains, for each triangle, the indices of its first triangle neighbors along its three edges. More precisely, we denote by $\mathcal{X}=\{x_{j}\}_{j=1,\dots,|\mathcal{X}|}$ the list of vertex coordinates $x_{j}\in\mathbb{R}^{2}$ where $|\mathcal{X}|$ is the total number of vertices, by $\mathcal{T}=\{T_{j}=(j_{1},j_{2},j_{3})\}_{j=1,\dots,|\mathcal{T}|}$ the list of index triplets defining the $j-$th triangle with vertices $x_{j_{1}},x_{j_{2}},x_{j_{3}}$ where $|\mathcal{T}|$ is the total number of triangles, and by $\mathcal{TN}=\{TN_{j}=(j_{1},j_{2},j_{3})\}_{j=1,\dots,|\mathcal{T}|}$ the list of index triplets corresponding to the three neighbors $T_{j_{1}},T_{j_{2}},T_{j_{3}},$ of the $j-$th triangle (for boundary elements one or two indices are set to zero if the corresponding neighbors are missing). Note that, in order to walk on the grid along the vertices, using $\mathcal{TN}$ is much cheaper than storing the first neighbors of each vertex, since the number of neighbors per vertex typically ranges from four up to eleven. Here we just store three indices per triangle, and we can use this connectivity list to move around a vertex, and accomplish the same task of finding vertex neighbors in constant time, by means of a custom function vertex_neighbors.
@@
This yields the following descent algorithm for the computation of the proposed node-restricted Hopf–Lax formula, presented in Algorithm $1$.
@@
7:             $k_{\text{old}}$leftarrow $\mathcal{K}^{d}_{j}$
@@
10:                 $k_{\text{new}}\,\leftarrow\,\displaystyle\arg\min_{k\in\mathcal{N}}\mathcal{F}(k)\,,\qquad\mathcal{F}(k)=\left\{v_{k}^{n}+\Delta tH^{*}\left(\frac{x_{j}-x_{k}}{\Delta t}\right)\right\}$
@@
14:             $v_{j,d}$leftarrow $\mathcal{F}({k_{\text{new}}})$
@@
16:         if $x_{j}$ on the boundary then
17:             $v_{j}^{n+1}\leftarrow\displaystyle\min{\{b_{j},\min_{d=1,\dots,4}v_{j,d}\}}$
18:         else
19:             $v_{j}^{n+1}\leftarrow\displaystyle\min_{d=1,\dots,4}v_{j,d}$
20:         end if
21:    end for
22:end for
@@
We remark that, under the assumption that the solution is convex, we can completely remove the initialization of walks based on different seeds. Actually, in such cases, $\mathcal{F}$ is also convex and has a unique global minimum. In this setting, inspired by [4], we can further optimize the process by starting the walk from the node that attained the minimum in the previous time step, and this leads to a significant speed-up in the overall computation.
We finally observe that, since we are interested in walks that strictly decrease $\mathcal{F}$, the function vertex_neighbors can be combined with an index-based comparison to restrict the evaluation of $\mathcal{F}$ only to the vertices that have not already been visited. This can clearly mitigate the overall computational cost of the algorithm in case the computation of $\mathcal{F}$ is particularly expensive.
The implementation of the proposed schemes for the stationary equation (7) on triangular grids parallels the one for the evolutive case presented in Section 2. In particular, we have to take into account the new terms related to the discount factor $\lambda$ and the source $f$ appearing in the right-hand side of (19). Moreover, for the value iteration scheme (24) we can readily apply Algorithm 1, just reinterpreting the number of time steps as the number of fixed-point iterations, and adding a stopping criterion based on some norm of the difference between two successive iterations. On the other hand, for both exact and modified policy iteration schemes, we have to embed our vertex neighbors minimum search in the policy improvement phase.

### v2

@@ 3 Construction of the scheme, stationary case
$u(x)=\inf_{\alpha(\cdot)}\left\{e^{-\lambda\Delta t}u(X_{\alpha}(\Delta t))+\int_{0}^{\Delta t}\Big(f(X_{\alpha}(s))+H^{*}(\alpha(s))\Big)e^{-\lambda s}ds\right\},$ (29)
where $\alpha\in L^{1}(\mathbb{R}^{+},\mathbb{R}^{d})$ and $X_{\alpha}(s)$ solves the ODE system (21).
In view of the $x$-dependence of $f$, (29) cannot be given by an exact Hopf–Lax type form; therefore, setting $y=X_{\alpha}(\Delta t)$, and $\alpha(s)\equiv(x-y)/\Delta t$, and assuming smoothness as usual, we approximate the integral in (29) as
$\int_{0}^{\Delta t}\Big(f(y(s;\alpha))+H^{*}(\alpha(s))\Big)e^{-\lambda s}ds=\Delta t\left(f(x)+H^{*}\left(\frac{x-y}{\Delta t}\right)\right)+O\left(\Delta t^{2}\right).$ (30)
@@
$v_{j}=\min_{k\in\mathbb{N}}\left\{e^{-\lambda\Delta t}v_{k}+\Delta tH^{*}\left(\frac{x_{j}-x_{k}}{\Delta t}\right)\right\}+\Delta tf(x_{j}).$ (32)
Equation (32) is a fixed-point equation, and it is easy to show that it has a contractive right-hand side (see [9]); this ensures that it has a unique solution $V$. Concerning the convergence of $V$ to the solution of (7), it is possible to retrace the arguments in [9], along with those of the time-dependent case above, to obtain consistency in the sense of [2], and a consistency estimate in the form
@@
While approximating the integral as in (30) suffices to obtain a consistent scheme for $\Delta x=o(\Delta t)$, the term $O(\Delta t)$ becomes a bottleneck for the consistency rate. An easy computation shows that the truncation error (33) is maximized taking $\Delta t\sim\Delta x^{2/3}$, and accordingly the consistency rate w.r.t. $\Delta x$ drops to 2/3. A way to increase the rate is to approximate the integral in the left-hand side of (30) by a trapezoidal formula,
$\int_{0}^{\Delta t}\Big(f(y(s;\alpha))+H^{*}(\alpha(s))\Big)e^{-\lambda s}ds$ $=$ $\frac{\Delta t}{2}\Big(f(x)+e^{-\lambda\Delta t}f(y)$ (34)
$+\big(1+e^{-\lambda\Delta t}\big)H^{*}\Big(\frac{x-y}{\Delta t}\Big)\Big)+O\left(\Delta t^{3}\right)$
(where we are still assuming $\alpha(s)\equiv(x-y)/\Delta t$ to be constant, and $f$ to be smooth enough), which results in the scheme
$v_{j}=\min_{k\in\mathbb{N}}\left\{e^{-\lambda\Delta t}v_{k}+\frac{\Delta t}{2}\left(\big(1+e^{-\lambda\Delta t}\big)H^{*}\left(\frac{x_{j}-x_{k}}{\Delta t}\right)+e^{-\lambda\Delta t}f(x_{k})\right)\right\}+\frac{\Delta t}{2}f(x_{j}).$ (35)
@@
Finally, we remark that, mixing the arguments of the time-dependent case with those of [9, Chapter 8] it is possible to prove error estimates of the form
$\|V^{n}-U^{n}\|_{\infty}\leq C_{1}\Delta t^{\gamma}+C_{2}\frac{\Delta x}{\Delta t},$
with $\gamma=1$ for the scheme (32), and $\gamma=2$ for the scheme (35).
3.1 Boundary conditions
Combining now the construction above with the techniques of the time-dependent case, we can modify the scheme for the case of weak Dirichlet conditions by adding a stopping cost on the boundary and considering the exit time as defined in (22). The form (32) of the scheme is then replaced by
$v_{j}=\min_{x_{k}\in\overline{\Omega}\atop\tau\in(0,\Delta t]}F_{j}(V,x_{k},\tau)$ (37)
where $V=(v_{1},\ldots,v_{N})$ and $F_{j}$ is defined as
[…]
The corresponding definition for the discretization (35) may be obtained with obvious changes. The optimal value $\tau_{jk}$ of $\tau$ for $x_{k}\in\partial\Omega$ may be computed following the guidelines of the time-dependent case, although in this case we do not expect an explicit solution, even for quadratic Hamiltonians. On the other hand, all the $\tau_{jk}$ can be computed once and for all at the start of the iterative solution of (37).
@@ 3.2 Value iteration
In the easiest approach, (32) (or (37)) may be solved by the so-called value iteration (VI), that is, via the iteration
$v_{j}^{(n+1)}=\min_{x_{k}\in\overline{\Omega}\atop\tau\in(0,\Delta t]}F_{j}\left(V^{(n)},x_{k},\tau\right)$ (39)
where the index $n$ should now be understood as an iteration number (an obvious modification applies for treating (35)). Since the right-hand side of (32) is a contraction, the iteration (39) converges to its unique fixed point. The contraction coefficient, however, is given by $e^{-\lambda\Delta t}$, and the convergence becomes very slow as the approximation is refined. This supports the use of faster solvers for the fixed-point equation, in particular the so-called Policy Iteration (PI), which will be briefly reviewed here. Note that, for shortness, we will outline the basic ideas on the case of an infinite grid, the introduction of boundary conditions being straightforward.
@@ 3.3 Exact policy iteration
    $\alpha^{(n+1)}\in\arg\min_{\alpha\in U^{N}}\left(B(\alpha)V^{(n)}-c(\alpha)\right),$
@@
$\min_{k_{j}\in\mathbb{N}}\left\{-v_{j}+e^{-\lambda\Delta t}v_{k_{j}}+\Delta t\left(H^{*}\left(\frac{x_{j}-x_{k_{j}}}{\Delta t}\right)+f(x_{j})\right)\right\}=0\quad(j=1,\ldots,N),$ (41)
where we understand the collection of all the indices $k_{j}$, i.e., the vector
${\bm{k}}=(k_{1},\ldots,k_{N})^{T}$
as the policy.
To interpret (41) in the form (40), it suffices to define the elements of the matrix $B(\bm{k})$ and the vector $c(\bm{k})$ as
$b_{jm}(\bm{k})=\begin{cases}1&\text{ if }m=j\neq k_{j}\\ -e^{-\lambda\Delta t}&\text{ if }m=k_{j}\neq j\\ 1-e^{-\lambda\Delta t}&\text{ if }m=j=k_{j}\\ 0&\text{ otherwise};\end{cases}\hskip 42.67912ptc_{j}(\bm{k})=\Delta t\left(H^{*}\left(\frac{x_{j}-x_{k_{j}}}{\Delta t}\right)+f(x_{j})\right).$ (42)
In this case, denoting by $k_{j}^{(n)}$ the $j$-th node index in the $n$-th policy $\bm{k}^{(n)}$, and changing all the signs, the policy evaluation phase is rewritten as the system
$v_{j}^{(n)}-e^{-\lambda\Delta t}v_{k_{j}^{(n)}}^{(n)}=\Delta t\left(H^{*}\left(\frac{x_{j}-x_{k_{j}^{(n)}}}{\Delta t}\right)+f(x_{j})\right)\quad(j=1,\ldots,N),$ (43)
while the policy improvement is computed as
$k_{j}^{(n+1)}\in\arg\min_{k_{j}\in\mathbb{N}}\left\{-v_{j}^{(n)}+e^{-\lambda\Delta t}v_{k_{j}}^{(n)}+\Delta t\left(H^{*}\left(\frac{x_{j}-x_{k_{j}}}{\Delta t}\right)+f(x_{j})\right)\right\}=0\quad(j=1,\ldots,N).$
With these definitions, assumptions (P1)(P2) of Theorem 3 are satisfied. In fact, the continuity of $B$ and $c$ is obvious once the set of node indices is naturally endowed with the discrete topology. As for (P2), we note that an equivalent definition of monotone matrix requires the off-diagonal elements to be nonpositive (which is clearly the case) and the eigenvalues to have nonnegative real part; this latter property is true, as a consequence of Gershgorin’s theorem. Note that monotonicity of the matrices $B$ implies their nonsingularity, and thus the unique solvability of the policy evaluation phase.
Clearly, the form (35) of the scheme can be treated by changing the definition of the vector $c(\bm{k})$ into
$c_{j}(\bm{k})=\frac{\Delta t}{2}\left(\big(1+e^{-\lambda\Delta t}\big)H^{*}\left(\frac{x_{j}-x_{k_{j}}}{\Delta t}\right)+f(x_{j})+e^{-\lambda\Delta t}f(x_{k_{j}})\right),$
@@ 3.4 Modified policy iteration
An adaptation of policy iteration to large sparse problems has been proposed as “modified policy iteration” (MPI) in [17], and has also become a classical tool. In this case, the policy evaluation phase is carried out in iterative form, which may be convenient in very large problems, whenever a direct solver would be unfeasible. In the specific case of (41), once given an initial guess $\bm{k}^{(0)}$ for the policy, the two nested cycles of MPI can be written as
@@
    $v_{j,m+1}^{(n)}=e^{-\lambda\Delta t}v_{k_{j}^{(n)},m}^{(n)}+\Delta t\left(H^{*}\left(\frac{x_{j}-x_{k_{j}^{(n)}}}{\Delta t}\right)+f(x_{j})\right)$
@@
    $k_{j}^{(n+1)}\in\arg\min_{k\in\mathbb{N}}\left\{-v_{j}^{(n)}+e^{-\lambda\Delta t}v_{k}^{(n)}+\Delta t\left(H^{*}\left(\frac{x_{j}-x_{k}}{\Delta t}\right)+f(x_{j})\right)\right\}.$
  3. 3.
    Stopping condition: if $\bm{k}^{(n+1)}\neq\bm{k}^{(n)}$ go to 1, else stop.
Here, we have denoted by $v_{j,m}^{(n)}$ the $j$-th element of the $m$-th iterate in the policy evaluation, by $v_{j}^{(n)}$ the corresponding final value of the iteration (i.e., the $j$-th element of $V^{(n)}$), and by $k_{j}^{(n)}$ the $j$-th element of the optimal policy at the $n$-th iteration. Note that, since the policy is defined, in this scheme, by the set of indices $k_{j}$ of the nodes achieving the minima for $F_{j}$, we can consider the policy to be convergent as soon as none of these indices is any longer updated.
Finally, note that all three solvers (Value Iteration, Exact Policy Iteration, Modified Policy Iteration) necessarily converge to the same solution.
@@ 4 Unstructured implementation
We provide some details for the actual implementation of the fully discrete scheme under consideration, on two-dimensional triangular grids. The aim is to compute the global minimum in (9) as efficiently as possible, in terms of both space and time complexity, by means of walks along the grid nodes that locally decrease the right hand side of our scheme. To this end, once a triangulation (typically, Delaunay) is given, we only assume to work with the minimal data structures describing the domain geometry, namely vertex and triangle lists, plus a connectivity list which contains, for each triangle, the indices of its first triangle neighbours along its three edges. More precisely, we denote by $\mathcal{X}=\{x_{j}\}_{j=1,\dots,|\mathcal{X}|}$ the list of vertex coordinates $x_{j}\in\mathbb{R}^{2}$ where $|\mathcal{X}|$ is the total number of vertices, by $\mathcal{T}=\{T_{j}=(j_{1},j_{2},j_{3})\}_{j=1,\dots,|\mathcal{T}|}$ the list of index triplets defining the $j$-th triangle with vertices $x_{j_{1}},x_{j_{2}},x_{j_{3}}$ where $|\mathcal{T}|$ is the total number of triangles, and by $\mathcal{TN}=\{TN_{j}=(j_{1},j_{2},j_{3})\}_{j=1,\dots,|\mathcal{T}|}$ the list of index triplets corresponding to the three neighbours $T_{j_{1}},T_{j_{2}},T_{j_{3}},$ of the $j$-th triangle (for boundary elements one or two indices are set to zero if the corresponding neighbours are missing). Note that, in order to walk on the grid along the vertices, using $\mathcal{TN}$ is much cheaper than storing the first neighbours of each vertex, since the number of neighbours per vertex typically ranges from four up to eleven. Here we just store three indices per triangle, and we can use this connectivity list to move around a vertex, and accomplish the same task of finding vertex neighbours in constant time, by means of a custom function vertex_neighbours.
@@
This yields the descent algorithm for the computation of the proposed node-restricted Hopf–Lax formula, presented in Algorithm 1.
@@
7:     $k_{\text{old}}$ $\leftarrow$ $\mathcal{K}^{d}_{j}$
@@
10:      $k_{\text{new}}\,\leftarrow\,\displaystyle\arg\min_{k\in\mathcal{N}}\mathcal{F}(k)\,,\qquad\mathcal{F}(k):=\min_{\tau\in(0,\Delta t]}F_{j}^{n}(x_{k},\tau)$    with $F_{j}^{n}$ given by (26)
@@
14:     $v_{j,d}$ $\leftarrow$ $\mathcal{F}({k_{\text{new}}})$
@@
16:    $v_{j}^{n+1}\leftarrow\displaystyle\min_{d=1,\dots,4}v_{j,d}$
17:  end for
18:end for
@@
We remark that the treatment of boundary conditions is embedded in the evaluation of $F_{j}^{n}$. Moreover, whenever the node $x_{k}$ lies on the boundary of the domain, the one-dimensional optimization of $F_{j}^{n}$ with respect to $\tau$ can be readily performed employing a standard root finder for its derivative, such as bisection or Newton method in case higher order derivatives can be explicitly computed.
As a rule of thumb for the choice of the constant $C$ in (44), note that, in order to handle an abrupt change in the minimum caused by the transit of a singularity, this constant should be of the order of (or slightly greater than) the maximum speed of propagation of the solution. For a Hamiltonian function of the form $H(p)$ used here, it is well known that the local speed of propagation at $x$ is $|\nabla H(Du(x))|_{2}$; on the other hand, by a zeroth-order Taylor expansion,
$|\nabla H(p)|_{2}$ $=$ $\left|\nabla H(0)+D^{2}H(\bar{p})p\right|_{2}$
$\leq$ $|\nabla H(0)|_{2}+\max\left|D^{2}H(\bar{p})\right|_{2}\max|p|_{2}.$
Since the gradient of the solution of (3) has a nonincreasing norm, $\max|p|_{2}$ may be bounded by $L_{0}$, the Lipschitz constant of $u_{0}$ in the Euclidean norm, while $\left|D^{2}H(\bar{p})\right|_{2}\leq M_{H}$ by the basic assumptions. Then, a reasonable choice of $C$ would be
$C\gtrsim|\nabla H(0)|_{2}+M_{H}L_{0}.$
For example, for the Hamiltonian $H(p)=\frac{1}{2}|p|^{2}$ which will be used in the numerical tests, we have $\nabla H(0)=0$ and $\left|D^{2}H(\bar{p})\right|_{2}\equiv 1$. In the stationary case, it is possible to replace $L_{0}$ with the Lipschitz constant of the time-discrete solution; for example, for the discrete formulation (31) the Lipschitz constant of $w$ is
$L_{w}=\frac{L_{f}}{\lambda},$
with $L_{f}$ denoting the Lipschitz constant of $f$.
Remark 2
Note that, assuming that the solution is convex, we can completely drop the initialization of walks based on different seeds. Actually, in this case, the right hand side in (23) is also convex and has a unique global minimum. Then, inspired by [6], we can further optimize the process by starting the walk from the node that attained the minimum in the previous time step, and this leads to a significant speed-up in the overall computation.
We finally observe that, since we are interested in walks that strictly decrease $F_{j}^{n}$, namely the right hand side of (25), the function vertex_neighbors can be combined with an index-based comparison to restrict the evaluation of $F_{j}^{n}$ only to the vertices that have not already been visited. This can clearly mitigate the overall computational cost of the algorithm in case the computation of $F_{j}^{n}$ is particularly expensive.
The implementation of the proposed schemes for the stationary equation (7) on triangular grids parallels the one for the evolutive case presented in Section 2. In particular, we have to take into account the new terms related to the discount factor $\lambda$ and the source $f$ appearing in the right-hand side of (32). Moreover, for the Value Iteration scheme (39) we can readily apply Algorithm 1, just reinterpreting the number of time steps as the number of fixed-point iterations, and adding a stopping criterion based on some norm of the difference between two successive iterations. On the other hand, for both Exact and Modified Policy Iteration schemes, we have to embed our vertex neighbours minimum search in the policy improvement phase.